• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kuchina

    Log entry time 08:00:31 on November 4,2006

    Entry number 188844

    keyword=owl shift summary

    Beam is off and on at the beginning of the shift.
    At about 2:30am David detected lost communication with RHRS
    3:15am problem solved
    Shortly after 4:30am beam was increased to 85uA
    At 6:20am MCC called, RF cavity problem

    Run summary
    1938 682k stopped:beam off
    1939 1M
    1940 282k stopped:lost communication with Right HRS
    1941 1M
    1942 1M
    1943 1M
    1944 137k stopped:beam off, MCC problem
    1945 1M
    1946 running